OpenStudy (anonymous):

What is the product of (2 – i)(4 + i)

OpenStudy (anonymous):

2-i=0 4+i=0 i=2 i=-4

OpenStudy (nubeer):

8 +2i-4i -i^2 8-2i +1 9-2i

hero (hero):

(2 – i)(4 + i) = 2(4 + i) -i(4 + i) = 8 + 2i - 4i + i^2 = 8 - 2i + 1 = 8 + 1 - 2i = 9 -2i